OpenStudy (annie96):

Find x. Select one: a. 46.2 b. 34.7 c. 23.1 d. 19.5

OpenStudy (kash_thesmartguy):

It asks for login, we don't have an account on there!

OpenStudy (anonymous):

|dw:1417565753369:dw

OpenStudy (kash_thesmartguy):

\[\frac{ 26 }{ 30 }=\frac{ x }{ 40 }\]\[40*0.87=34.8\] x=34.8

OpenStudy (kash_thesmartguy):

Or, 34 .7 so B.

OpenStudy (kash_thesmartguy):

You get? @Annie96

OpenStudy (annie96):

Yes thank you @Kash_TheSmartGuy

OpenStudy (kash_thesmartguy):

np :)